• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 13:07:49 on October21,2002

    Entry number 88633

    keyword=Moved to Lh2-15 cm target position

    Just moved the target from EMPTY to LH2_15cm (13:05).
    Target was in 12C at teh start of shift and moved to LH2 after
    the current scan completed for 12C (around 9:30). Then moved
    to EMPTY for beam study before the controlled access.